. A lake is 10 m deep and 1 km by 1 km in area. All the water goes over a waterfall of 300 m high. How much electrical energy is produced at 100% efficiency from this energy supply?
1
Expert's answer
2020-08-21T10:44:16-0400

The mass of water in a lake

"m=\rho V\\
=1000\:\rm kg/m^3\times 10\: m\times 1000\: m\times 1000\: m=10^{10}\: kg"

The amount of energy

"E=mgh\\
=10^{10}\:\rm kg\times 9.8\: N/kg\times 300\: m=2.94\times 10^{13}\:\rm J"